OpenStudy (henryarias14):

CRITICAL POINTS This is the problem x^2 / (x^2+3) I found the first derivative 6x/(x^2+3)^2 But I don't know what to do next to find the zeroes. I know I have to equal the derivative to 0 and I did but then I got lost. I'd appreciate any help, thank you

OpenStudy (henryarias14):

On the closed interval [-1,1]

OpenStudy (henryarias14):

Can you help me out?

zepdrix (zepdrix):

Derivative looks good :)\[\large\rm 0=\frac{6x}{(x^2+3)^2}\]Multiplying through by our denominator gives us,\[\large\rm 0=6x\]

OpenStudy (henryarias14):

How did you get 6x? Sorry I need to refresh my algebra

zepdrix (zepdrix):

Multiplying both sides by \(\large\rm (x^2+3)^2\) Maybe a good shortcut to remember is just to `set your numerator equal to zero`.

zepdrix (zepdrix):

Now, something you should realize though... We also must set our denominator equal to zero. \(\large\rm 0=(x^2+3)^2\) This corresponds to a `different type of critical point`, but it's still a critical point. It doesn't refer to locations of horizontal tangency, but instead locations of vertical tangency.

OpenStudy (henryarias14):

Ohhhh since the zero is on the other side it cancels out. Yeah I need to remember setting it equal to zero, Thanks!! So the critical point would be x=-6

zepdrix (zepdrix):

No no :O woops! 0=6x, the 6 is `multiplying` x. So you have to `divide`.

OpenStudy (henryarias14):

Right. So x=0 again I need to remember setting it equal to zero

OpenStudy (henryarias14):

Now, if I wanted to find the absolute extrema I need to plug in 0 into the derivative or the original equation?

OpenStudy (henryarias14):

Had to leave class early so I kind of had to learn it alone. It sucks

zepdrix (zepdrix):

Critical points refer to locations of horizontal tangency. `The slope of the function at these critical points is zero`. So if you plugged your critical point back into the derivative, it would give you zero. That's what we figured out by setting the derivative equal to zero. So that won't do us any good :) Plug it back into the original function get an exact location (coordinate pair) of your extrema.

OpenStudy (henryarias14):

If I plug 0 into the original equation I get (0,0)

zepdrix (zepdrix):

So this min or max point (whatever it may be) goes through the origin? Yup looks correct! :)

OpenStudy (henryarias14):

Don't I have to plug in the intervals too? It's set on a closed interval [-1,1] don't I have to plug them too?

OpenStudy (henryarias14):

Trying to find the maxima

zepdrix (zepdrix):

Oh oh oh oh.. yes sorry >.< When you're given a closed interval, you check for critical points, then you take all of the important points that you have: critical points and end points, and plug them into the original function, and compare their outputs.

OpenStudy (henryarias14):

Mmm but I get (-1,1/4) and (1,1/4) which one will be the maxima? I mean they're both the same

zepdrix (zepdrix):

Maximum: (-1, 1/4) and (1, 1/4) Minimum: (0,0) Yes :) good job! We list critical points by their x-coordinate, but extrema as a coordinate pair.

OpenStudy (henryarias14):

Oh I see. Thanks so much! I have another problem which is just finding a derivative that I wouldn't find if you have time?

OpenStudy (henryarias14):

Couldn't*

OpenStudy (henryarias14):

It's f(x)= x(sqr 4-x)

OpenStudy (henryarias14):

I'm on a phone so I can't use the symbols

zepdrix (zepdrix):

\[\large\rm f(x)=x\cdot\sqrt{4-x}\]Hmm any ideas how to start this one? We have two things multiplying together, anything clicking? XD

OpenStudy (henryarias14):

Lol yeah I mean I got (1)(sqr 4-x) +

OpenStudy (henryarias14):

And then that part I didn't get lol I know it's product rule and.... chain?

OpenStudy (henryarias14):

I think it's called a composite derivative or something like that

zepdrix (zepdrix):

Early on, it might be a good idea to spend an extra step "setting up" your product rule. If you're comfortable skipping that, it's fine.\[\large\rm f'(x)=(x)'\sqrt{4-x}+x(\sqrt{4-x})'\]

zepdrix (zepdrix):

\[\large\rm f'(x)=(1)\sqrt{4-x}+x(\sqrt{4-x})'\]

OpenStudy (henryarias14):

Yeah Ima confortable skipping it but I got all messed up trying to find the derivative of the square inside the product

OpenStudy (henryarias14):

Im*

zepdrix (zepdrix):

You can convert your root to a 1/2 power if you like, but honestly, the square root comes up soooo oftne, it's really worth your time to memorize the square root derivative.

OpenStudy (henryarias14):

I see

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt x}\]Derivative of square root is one over two square roots.

zepdrix (zepdrix):

So yes, you'll have to apply the chain rule as extra step since our inner function is `more than just x`,\[\large\rm \frac{d}{dx}\sqrt{4-x}\quad=\quad \frac{1}{2\sqrt{4-x}}\cdot (4-x)'\]

OpenStudy (henryarias14):

So how would you set it up? The chain rule and the product rule together for this problem

OpenStudy (henryarias14):

The chain part of the problem would be: (x) • 1/2(sqr 4-x) • (-1) ...... (-1) coming from the derivative of 4-x

zepdrix (zepdrix):

Well, you chain every time you take a derivative. That let's us know that product rule will be applied first, ya? Because the "setting up" of our product rule isn't actually the process of taking a derivative. Yes that looks good. Lemme recap just in case it will help...

OpenStudy (henryarias14):

Alright. I wish I could send pictures through here

zepdrix (zepdrix):

Oh you can't see the LaTex formatted? :( Ugh, I was gonna do it in color >.< heh

OpenStudy (henryarias14):

What do you mean? I pretty much can see everything I just can't send it

zepdrix (zepdrix):

Haha I misread your last line, read it as "I wish I could `see` pictures through here" XD lol

OpenStudy (henryarias14):

Lol I figured

zepdrix (zepdrix):

So I like to use this color theme: Blue are derivatives that need to be taken, Orange is when we finish them. So we set up our product rule, \[\large\rm f'(x)=\color{royalblue}{(x)'}\sqrt{4-x}+x\color{royalblue}{(\sqrt{4-x})'}\]Take our first derivative,\[\large\rm f'(x)=\color{orangered}{(1)}\sqrt{4-x}+x\color{royalblue}{(\sqrt{4-x})'}\]Take the other one, creating a chain,\[\large\rm f'(x)=\color{orangered}{(1)}\sqrt{4-x}+x\color{orangered}{\frac{1}{2\sqrt{4-x}}}\color{royalblue}{(4-x)'}\]Finishing up,\[\large\rm f'(x)=\color{orangered}{(1)}\sqrt{4-x}+x\color{orangered}{\frac{1}{2\sqrt{4-x}}}\color{orangered}{(-1)}\]

zepdrix (zepdrix):

And then maybe getting a common denominator from there. But whatever.

zepdrix (zepdrix):

So one tiny thing to notice is that the chain is only multiplying the second term, so we don't end up distributing that -1 to each term or anything like that.

OpenStudy (henryarias14):

Yeah I noticed. Now how would the next step look after multiplying the 1, -1, and x? Cause I did it on my own but i got one whole fraction

OpenStudy (henryarias14):

(Sqr 4-x) - x all over (sqr 4-x)

zepdrix (zepdrix):

\[\large\rm f'(x)=\sqrt{4-x}-\frac{x}{2\sqrt{4-x}}\]Hmm I guess that first term needs a 2sqrt(4-x) to have the same denominator, ya?\[\large\rm f'(x)=\left(\frac{2\sqrt{4-x}}{2\sqrt{4-x}}\right)\sqrt{4-x}-\frac{x}{2\sqrt{4-x}}\]Giving us,\[\large\rm f'(x)=\frac{2(4-x)}{2\sqrt{4-x}}-\frac{x}{2\sqrt{4-x}}\]combining the fractions,\[\large\rm f'(x)=\frac{2(4-x)-x}{2\sqrt{4-x}}\]Looks like the square root disappears from the numerator, ya?

OpenStudy (henryarias14):

Mmmm yeah it disappears but we still have it in the denominator

zepdrix (zepdrix):

So ya I guess you could go a step further, distribute the 2, and then combine like-terms.\[\large\rm f'(x)=\frac{8-3x}{2\sqrt{4-x}}\]

OpenStudy (henryarias14):

Oh what I did was I put everything under the same numerator but It was wrong

OpenStudy (henryarias14):

That would be the final form of the derivative?

OpenStudy (henryarias14):

Oh I see now

OpenStudy (henryarias14):

CRitical point would be 8/3? Lol

zepdrix (zepdrix):

Ya, that's a good looking final answer. The earlier one was fine also though,\[\large\rm f'(x)=\sqrt{4-x}-\frac{x}{2\sqrt{4-x}}\]Just depends how far your teacher wants you to simplify things.

OpenStudy (henryarias14):

Yeah he wants us to go all the way when it comes to simplifying

zepdrix (zepdrix):

You might have missed the subtle point I made earlier, Yes, 8/3 would end up being a critical value for the function. But we also have to set the denominator equal to zero to find other critical values.\[\large\rm 0=2\sqrt{4-x}\]

zepdrix (zepdrix):

The reason we didn't worry about this in the last problem is because: The value we would have ended up with was `not in the domain of our original function`. The function and derivative had the same factor in the denominator, so the solution wasn't allowed. That's not the case for this problem though, you'll need to solve this equation to find the other critical point.

OpenStudy (henryarias14):

Yeah but I thought if we multiplied both sides by the denominator, it'll end up canceling out cause the 0 is on the other side. Then we end up having 8-3x=0

zepdrix (zepdrix):

It's a different type of critical point :) It's a weird one.

OpenStudy (henryarias14):

Well the problem added that x<3, they didn't give me an interval just asking for critical points

OpenStudy (henryarias14):

Idk if the numerator fits x<3

OpenStudy (henryarias14):

Denominator***

zepdrix (zepdrix):

So you found this critical point. The slope of the line tangent to the curve is zero (horizontal).